# Generalized Random Forests using Fixed-Point Trees

### v3

@@ Abstract
We propose a computationally efficient alternative to generalized random forests (GRFs) for estimating heterogeneous effects in large dimensions. While GRFs rely on a gradient-based splitting criterion, which in large dimensions is computationally expensive and unstable, our method introduces a fixed-point approximation that eliminates the need for Jacobian estimation. This gradient-free approach preserves GRF’s theoretical guarantees of consistency and asymptotic normality while significantly improving computational efficiency. We demonstrate that our method achieves multiple times the speed over standard GRFs without compromising statistical accuracy. Experiments, on both simulated and real-world data, validate our approach. Our findings suggest that the proposed method is a scalable alternative for localized effect estimation in machine learning and causal inference applications.
@@ 1 Introduction
Generalized random forests (GRFs) (Athey et al., 2019; Wager & Athey, 2018) have emerged as a powerful tool for such tasks, leveraging adaptive partitioning with problem-specific moment conditions instead of standard loss-based splits. Unlike local linear (Fan et al., 1995; Fan & Gijbels, 1996; Friedberg et al., 2020) or kernel-based models (Staniswalis, 1989; Severini & Staniswalis, 1994; Lewbel, 2007; Speckman, 1988; Robinson, 1988), which suffer from the curse of dimensionality (Robins & Ritov, 1997), GRFs offer a scalable alternative.
@@ 2 Background and Related Work
$0=\mathbb{E}_{O|X}\left[\psi_{\theta^{*}(x),\nu^{*}(x)}(O_{i})\,\middle|\,X_{i}=x\right],$ (1)
for all $x\in\mathcal{X}$, where $\psi$ is a score function that identifies the true $(\theta^{*}(x),\nu^{*}(x))$ as the root of (1), and $\nu^{*}(x)$ is an optional nuisance function. GRF can be understood from a nearest-neighbour perspective as approximating $\theta^{*}(x)$ through a locally parametric $\theta^{*}$ within small neighborhoods of test point $x$. Suppose $L(x)\subset\{X_{i}\}_{i=1}^{n}$ is a subset of training observations of the covariates found in a region around $x\in\mathcal{X}$ over which $\theta^{*}(x)$ can be well-approximated by a local parameter. Observations $X_{i}\in L(x)$ serve as local representatives for $x$ in estimating $\theta^{*}(x)$ such that, given sufficiently many training samples in a small enough neighbourhood of $x$, an empirical version of (1) over $X_{i}\in L(x)$ defines an estimator $\hat{\theta}_{L(x)}$ that approaches $\theta^{*}(x)$,
@@
In GRF, the set of local representatives $L(x)$ is determined by tree-based partitions which divide the input space into disjoint regions, or leaves. The training samples $X_{i}$ that fall in the same leaf as $x$ form the subset $L(x)$. However, single trees are known to have high variance with respect to small changes in the training data (Amit & Geman, 1997; Breiman, 1996, 2001; Dietterich, 2000), leading to estimates (2) that do not generalize well values of $x$ that are not part of the training set. GRF improves its estimates by leveraging an estimating function that averages many estimating functions of the form (2). Specifically, let $L_{b}(x)$ denote the set of training covariates that fall in the same leaf as $x$, identified by a tree trained on an independent subsample of the data, indexed by $b=1,\ldots,B$. The GRF estimator is obtained by aggregating the individual estimating functions (2) across a forest of $B$ independently trained trees, i.e. the solution to the following forest-averaged estimating equation:
@@
Our contribution improves the computational cost of Stage I by introducing a more efficient procedure to train the trees. Training the forest is the most resource-intensive step of GRF, and the cost of each split in the existing approach scales quadratically with dimension of $\theta^{*}(x)$. We adopt a gradient-free splitting mechanism and significantly reduce both the time and memory demands of Stage I. Crucially, solving Stage II with weights $\alpha_{i}(x)$ following our streamlined Stage I produces an estimator $\hat{\theta}(x)$ that preserves the finite-sample performance and asymptotic guarantees of GRF.
@@ 3.1 The target tree-splitting criterion for Stage I
In GRF, the goal of Stage I is to use recursive tree-based splits of the training data to induce a partition over the input space. Each split starts with a parent node $P\subset\mathcal{X}$ and results in child nodes $C_{1},C_{2}\subset\mathcal{X}$, defined by a binary, axis-aligned splitting rule of the form $C_{1}=\{X_{i}\,:\,X_{i,j}\leq s\}$ and $C_{2}=\{X_{i}\,:\,X_{i,j}>s\}$, where $j$ denotes a candidate feature and $s$ denotes the corresponding level used to perform a split. For a parent $P$ and any child nodes $C_{1},C_{2}$ of $P$, let $(\hat{\theta}_{P},\hat{\nu}_{P})$ and $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$ denote local solutions analogous to (2) defined over the samples in $P$ and $C_{j}$, respectively:
@@
for $j=1,2$. A strategy to split $P$ into two subsets of greater homogeneity with respect to $\theta^{*}(\cdot)$ is as follows: Find child nodes $C_{1}$ and $C_{2}$ such that the total deviation between the local solutions $\hat{\theta}_{C_{j}}$ and the target $\theta^{*}(X)$ is minimized, conditional to $X\in C_{j}$, $j=1,2$. A natural measure of deviation is the squared-error loss,
$\textnormal{err}(C_{1},C_{2})\coloneqq$ $\sum_{j=1,2}\mathbb{P}\left(X\in C_{j}\middle|X\in P\right)$
@@
such that the resulting split $\{C_{1},C_{2}\}$ corresponds to least-squares optimal solutions $\hat{\theta}_{C_{1}}$ and $\hat{\theta}_{C_{2}}$. However, $\textnormal{err}(C_{1},C_{2})$ is intractable since $\theta^{*}(\cdot)$ is unknown. GRF considers a criterion that measures heterogeneity across a pair of local solutions over a candidate split
$\Delta(C_{1},C_{2})\coloneqq\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}\right\rVert^{2},$
@@
Although the criterion $\Delta(C_{1},C_{2})$ is computable, evaluating it is very computationally expensive since it requires solving (7) to obtain $\hat{\theta}_{C_{1}},\hat{\theta}_{C_{j}}$ for all possible splits of $P$, and closed-form solutions for $\hat{\theta}_{C_{j}}$ are generally not available except in special cases of $\psi$. Instead, GRF approximates the target $\Delta$-criterion based on a criterion of the form
@@
where $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$ denotes a gradient-based approximation of $\hat{\theta}_{C_{j}}$. Specifically, $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$ is a first-order approximation interpreted as the result of taking a gradient step away from the parent estimate in the direction towards the true local solution $\hat{\theta}_{C_{j}}$:
@@
where $(\hat{\theta}_{P},\hat{\nu}_{P})$ is the local solution over the parent, $A_{P}$ is any consistent estimator of the local Jacobian matrix $\nabla_{(\theta,\nu)}\mathbb{E}[\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})\,|\,X_{i}\in P]$, and $\xi^{\top}$ can be thought of as a term that selects a $\theta$-subvector from a $(\theta,\nu)$-vector, e.g. if $\theta\in\mathbb{R}^{K}$ and $\nu\in\mathbb{R}$, then $\xi^{\top}$ such that $\theta=\xi^{\top}(\theta,\nu)^{\top}$ is the rectangular diagonal matrix $\xi^{\top}=\left[\begin{matrix}\mathbb{I}_{K}&{\bf 0}\end{matrix}\right]$. When the scoring function $\psi$ is continuously differentiable in $(\theta,\nu)$, the Jacobian estimator $A_{P}$ can be computed as
@@ 3.2 Limitations of gradient-based approximation
The use of the Jacobian estimator $A_{P}$ in (9) introduces considerable computational challenges. First, each parent node $P$ in every tree of the forest requires a distinct $A_{P}$ matrix, which imposes a significant computational burden when explicitly caculating $A_{P}^{-1}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$ to determine $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$. Second, if the local Jacobian $\nabla_{(\theta,\nu)}\mathbb{E}[\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})\,|\,X_{i}\in P]$ is ill-conditioned, then the resulting $A_{P}$ estimator may be nearly singular. This instability can lead to highly variable gradient-based approximations $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$ and highly variable splits of $P$. For example, consider the following varying-coefficient model for an outcome $Y_{i}$ given regressors $W_{i}=(W_{i,1},\ldots,W_{i,K})^{\top}$ in the presence of mediating auxiliary covariates $X_{i}$:
$\mathbb{E}[Y_{i}\,|\,X_{i}=x]=\nu^{*}(x)+W_{i}^{\top}\theta^{*}(x),$ (11)
@@
$\psi_{\theta,\nu}(Y_{i},W_{i})\coloneqq\begin{bmatrix}Y_{i}-W_{i}^{\top}\theta-\nu\\ (Y_{i}-W_{i}^{\top}\theta-\nu)W_{i}\end{bmatrix}.$
Consequently, the local Jacobian estimator $A_{P}$ defined in (3.1) is
$A_{P}=$ $\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}\nabla_{(\nu,\theta)}\psi_{\theta,\nu}(Y_{i},W_{i})$
$=$ $-\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}\begin{bmatrix}1&W_{i}^{\top}\\ W_{i}&W_{i}W_{i}^{\top}\end{bmatrix}.$ (12)
When the regressors are highly correlated, the summation over the $W_{i}W_{i}^{\top}$ block of the $A_{P}$ matrix leads to nearly singular values of $A_{P}$, resulting in an unstable matrix inverse $A_{P}^{-1}$, and therefore unstable values of $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$ and unstable splits. This issue becomes more pronounced as the number of samples $n_{P}$ in a parent node $P$ decreases, particularly at deeper levels of the tree. These challenges highlights the limitations of relying on $A_{P}$ in such scenarios.
@@ 3.3 Fixed-point approximation
To address the limitations of gradient-based approximations, we propose a gradient-free approach based on fixed-point iteration. Let $\Psi_{C_{j}}(\theta,\nu)\coloneqq\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\theta,\nu}(O_{i})$ denote the empirical estimating function for the local solution $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$ defined in (7). This solution is obtained by solving the optimization problem:
$(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})\in\operatorname*{arg\,min}_{\theta,\nu}\left\lVert\Psi_{C_{j}}(\theta,\nu)\right\rVert,$ (14)
for $j=1,2$. Under mild regularity conditions, $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$ is a $Z$-estimator that is identified by the estimating equation $\Psi_{C_{j}}(\theta,\nu)={\bf 0}$. Reformulating this equation as a fixed-point problem, we write:
$(\theta,\nu)=(\theta,\nu)-\eta\Psi_{C_{j}}(\theta,\nu)\eqqcolon f(\theta,\nu),$ (15)
for any $\eta>0$. Specifically, a necessary and sufficient condition for $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$ to be a solution of (14) is characterized by the following fixed-point problem $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})=f(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$, where $f$ is as defined in (15). Iterative fixed-point iteration methods (Picard, 1890; Lindelöf, 1894; Banach, 1922; Ryu & Boyd, 2016) solve such problems by considering updates of the form
@@
This inspires us to approximate the true local solution $\hat{\theta}_{C_{j}}$ using a single fixed-point iteration taken from the parent solution $\hat{\theta}_{P}$:
@@
$=$ $\hat{\theta}_{P}-\xi^{\top}\frac{\eta}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}),$ (17)
where $\xi^{\top}$ excludes the nuisance terms from the update in (16). We interpret $\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}$ as an approximation of $\hat{\theta}_{C_{j}}$ obtained by taking a step from $\hat{\theta}_{P}$ in a direction that reduces the magnitude of the local estimating function $\Psi_{C_{j}}$. Notably, the approximations $\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ does not involve the $A_{P}$ matrix, relying only on evaluations of the parent scores $\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$. The corresponding splitting criterion, which uses the fixed-point approximations $\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ as substitutes for $\hat{\theta}_{C_{j}}$ in the target $\Delta$-criterion, is given by
@@
Revisiting the varying coefficient example from Section 3.2, we see that splits based on fixed-point approximations $\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ are significantly more stable those based on $\tilde{\theta}_{C_{j}}^{\texttt{grad}}$. Figure 1 illustrates that splits which maximize the fixed-point criterion $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ are significantly more robust to ill-conditioned local Jacobian matrices $\nabla_{(\theta,\nu)}\mathbb{E}[\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})\,|\,X_{i}\in P]$, as is the case for highly correlated regressors in the varying coefficient model (13), leading to highly stable splits.

### v4

@@ Abstract
We propose a computationally efficient alternative to generalized random forests (GRFs) for estimating heterogeneous effects in large dimensions. While GRFs rely on a gradient-based splitting criterion, which in large dimensions is computationally expensive and unstable, our method introduces a fixed-point approximation that eliminates the need for Jacobian estimation. This gradient-free approach preserves GRF’s theoretical guarantees of consistency and asymptotic normality while significantly improving computational efficiency. We demonstrate that our method achieves a speedup of multiple times over standard GRFs without compromising statistical accuracy. Experiments on both simulated and real-world data validate our approach. Our findings suggest that the proposed method is a scalable alternative for localized effect estimation in machine learning and causal inference applications.
@@ 1 Introduction
Generalized random forests (GRFs) (Athey et al., 2019; Wager & Athey, 2018) have emerged as a powerful tool for such tasks, leveraging adaptive partitioning with problem-specific moment conditions instead of standard loss-based splits. GRFs apply broadly to a wide range of important statistical models – local linear regression (Friedberg et al., 2020), survival analysis and missing data problems (Cui et al., 2023), nonparametric quantile regression, heterogeneous treatment effect estimation, and nonlinear instrumental variables regression (Athey & Imbens, 2016; Athey et al., 2019). Unlike local linear models (Fan et al., 1995; Fan & Gijbels, 1996; Friedberg et al., 2020) or kernel-based models (Staniswalis, 1989; Severini & Staniswalis, 1994; Lewbel, 2007; Speckman, 1988; Robinson, 1988) which suffer from the curse of dimensionality (Robins & Ritov, 1997), the tree-based approach of GRF offers a more scalable solution.
@@ 2 Background and Related Work
$0=\mathbb{E}_{O|X}\left[\psi_{\theta^{*}(x),\nu^{*}(x)}(O)\mid X=x\right],$ (1)
for all $x\in\mathcal{X}$, where $\psi$ is a score function that identifies the true $(\theta^{*}(x),\nu^{*}(x))$ as the root of (1), and $\nu^{*}(x)$ is an optional nuisance function. GRF can be understood from a nearest-neighbor perspective as approximating $\theta^{*}(x)$ through a locally parametric $\theta^{*}$ within small neighborhoods of test point $x$. Suppose $L(x)\subset\{X_{i}\}_{i=1}^{n}$ is a subset of training observations of the covariates found in a region around $x\in\mathcal{X}$ over which $\theta^{*}(x)$ can be well-approximated by a local parameter. Observations $X_{i}\in L(x)$ serve as local representatives for $x$ in estimating $\theta^{*}(x)$ such that, given sufficiently many training samples in a small enough neighborhood of $x$, an empirical version of (1) over $X_{i}\in L(x)$ defines an estimator $\hat{\theta}_{L(x)}$ that approaches $\theta^{*}(x)$,
@@
In GRF, the set of local representatives $L(x)$ is determined by tree-based partitions which divide the input space into disjoint regions, or leaves. The training samples $X_{i}$ that fall in the same leaf as $x$ form the subset $L(x)$. However, single trees are known to have high variance with respect to small changes in the training data (Amit & Geman, 1997; Breiman, 1996, 2001; Dietterich, 2000), leading to estimates (2) that do not generalize well to values of $x$ that are not part of the training set. GRF improves its estimates by leveraging an estimating function that averages many estimating functions of the form (2). Specifically, let $L_{b}(x)$ denote the set of training covariates that fall in the same leaf as $x$, identified by a tree trained on an independent subsample of the data, indexed by $b=1,\ldots,B$. The GRF estimator is obtained by aggregating the individual estimating functions (2) across a forest of $B$ independently trained trees, i.e. the solution to the following forest-averaged estimating equation:
@@
Our contribution improves the computational cost of Stage I by introducing a more efficient procedure to train the trees. Training the forest is the most resource-intensive step of GRF, and the cost of each split in the existing approach scales quadratically with the dimension of $\theta^{*}(x)$. We adopt a gradient-free splitting mechanism and significantly reduce both the time and memory demands of Stage I. Crucially, solving Stage II with weights $\alpha_{i}(x)$ following our streamlined Stage I produces an estimator $\hat{\theta}(x)$ that preserves the finite-sample performance and asymptotic guarantees of GRF.
@@ 3.1 The target tree-splitting criterion for Stage I
In GRF, the goal of Stage I is to use recursive tree-based splits of the training data to induce a partition over the input space. Each split starts with a parent node $P\subset\mathcal{X}$ and results in child nodes $C_{1},C_{2}\subset\mathcal{X}$, defined by a binary, axis-aligned splitting rule of the form $C_{1}=\{X_{i}\,:\,X_{i,\ell}\leq t\}$ and $C_{2}=\{X_{i}\,:\,X_{i,\ell}>t\}$, where $\ell$ denotes a candidate splitting feature/axis and $t\in\mathbb{R}$ the splitting threshold. For a parent $P$ and any child nodes $C_{1},C_{2}$ of $P$, let $(\hat{\theta}_{P},\hat{\nu}_{P})$ and $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$ denote local solutions analogous to (2) defined over the samples in $P$ and $C_{j}$, respectively:
@@
for $j=1,2$. A strategy to split $P$ into two subsets of greater homogeneity with respect to $\theta^{*}(\cdot)$ is as follows: Find child nodes $C_{1}$ and $C_{2}$ such that the total deviation between the local solutions $\hat{\theta}_{C_{j}}$ and the target $\theta^{*}(X)$ is minimized, conditional on $X\in C_{j}$, $j=1,2$. A natural measure of deviation is the squared-error loss,
$\textnormal{err}(C_{1},C_{2})\coloneqq$ $\sum_{j=1,2}\mathbb{P}\left(X\in C_{j}\mid X\in P\right)$
@@
such that the resulting split $(C_{1},C_{2})$ corresponds to least-squares optimal solutions $\hat{\theta}_{C_{1}}$ and $\hat{\theta}_{C_{2}}$. However, $\textnormal{err}(C_{1},C_{2})$ is intractable since $\theta^{*}(\cdot)$ is unknown. GRF considers a criterion that measures heterogeneity across a pair of local solutions over a candidate split
$\Delta(C_{1},C_{2})\coloneqq\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}\right\rVert^{2},$ (8)
@@
Although the criterion $\Delta(C_{1},C_{2})$ is computable, evaluating it is very computationally expensive since it requires solving (7) to obtain $\hat{\theta}_{C_{1}},\hat{\theta}_{C_{2}}$ for all possible splits of $P$, and closed-form solutions for $\hat{\theta}_{C_{j}}$ are generally not available except in special cases of $\psi$. Instead, GRF approximates the target $\Delta$-criterion based on a criterion of the form
@@
where $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$ denotes a gradient-based approximation of $\hat{\theta}_{C_{j}}$. Specifically, $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$ is a first-order approximation interpreted as the result of taking a gradient step away from the parent estimate in the direction towards the true child solution $\hat{\theta}_{C_{j}}$:
@@
where $(\hat{\theta}_{P},\hat{\nu}_{P})$ is the local solution over the parent, $A_{P}$ is any consistent estimator of the local Jacobian matrix $\nabla_{(\theta,\nu)}\mathbb{E}[\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})\mid X_{i}\in P]$, and $\xi^{\top}$ can be thought of as a term that selects a $\theta$-subvector from a $(\theta,\nu)$-vector, e.g. if $\theta\in\mathbb{R}^{K}$ and $\nu\in\mathbb{R}$, then $\xi^{\top}$ such that $\theta=\xi^{\top}(\theta,\nu)^{\top}$ is the rectangular diagonal matrix $\xi^{\top}=\left[\begin{matrix}\mathbb{I}_{K}&{\bf 0}\end{matrix}\right]$. When the scoring function $\psi$ is continuously differentiable in $(\theta,\nu)$, the Jacobian estimator $A_{P}$ can be computed as
@@ 3.2 Limitations of gradient-based approximation
The use of the Jacobian estimator $A_{P}$ in (10) introduces considerable computational challenges. First, each parent node $P$ in every tree of the forest requires a distinct $A_{P}$ matrix, which imposes a significant computational burden when explicitly calculating $A_{P}^{-1}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$ to determine $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$. Second, if the local Jacobian $\nabla_{(\theta,\nu)}\mathbb{E}[\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})\mid X_{i}\in P]$ is ill-conditioned, then the resulting $A_{P}$ estimator may be nearly singular. This instability can lead to highly variable gradient-based approximations $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$ and highly variable splits of $P$. For example, consider the following varying-coefficient model for an outcome $Y_{i}$ given regressors $W_{i}=(W_{i,1},\ldots,W_{i,K})^{\top}$ in the presence of mediating auxiliary covariates $X_{i}$:
$\mathbb{E}[Y_{i}\mid X_{i}=x]=\nu^{*}(x)+W_{i}^{\top}\theta^{*}(x),$ (12)
@@
$\psi_{\theta,\nu}(Y_{i},W_{i})\coloneqq\begin{bmatrix}(Y_{i}-W_{i}^{\top}\theta-\nu)\cdot W_{i}\\ Y_{i}-W_{i}^{\top}\theta-\nu\end{bmatrix}.$
Consequently, the corresponding local Jacobian estimator is
$A_{P}=$ $\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}\nabla_{(\theta,\nu)}\psi_{\theta,\nu}(Y_{i},W_{i})$
$=$ $-\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}\begin{bmatrix}W_{i}W_{i}^{\top}&W_{i}^{\top}\\ W_{i}&1\end{bmatrix}.$ (13)
When the regressors are highly correlated, the summation over the $W_{i}W_{i}^{\top}$ block of the $A_{P}$ matrix leads to nearly singular values of $A_{P}$, resulting in an unstable matrix inverse $A_{P}^{-1}$, and therefore unstable values of $\tilde{\theta}^{\texttt{grad}}_{C_{j}}$ and unstable splits. This issue becomes more pronounced as the number of parent samples $n_{P}$ decreases, as is the case at deeper levels of the tree. These challenges highlight the limitations of relying on $A_{P}$ as part of an approximation for the child solutions $\hat{\theta}_{C_{j}}$.
@@ 3.3 Fixed-point approximation
To address the limitations of gradient-based approximations, we propose a gradient-free approach based on the form of a single fixed-point iteration. Let $\Psi_{C_{j}}(\theta,\nu)\coloneqq\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\theta,\nu}(O_{i})$ denote the empirical estimating function for the child solution $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$ such that (7) is equivalently written as:
$(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})\in\operatorname*{arg\,min}_{\theta,\nu}\left\lVert\Psi_{C_{j}}(\theta,\nu)\right\rVert,\quad j=1,2.$ (15)
Under mild regularity conditions, $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$ is a $Z$-estimator that solves the estimating equation $\Psi_{C_{j}}(\theta,\nu)={\bf 0}$. Reformulating this equation as a fixed-point problem, we write:
$(\theta,\nu)=\underbrace{(\theta,\nu)-\eta\Psi_{C_{j}}(\theta,\nu)}_{\eqqcolon f(\theta,\nu)},\quad\eta>0.$ (16)
A necessary and sufficient condition for $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$ to be a solution of (15) is characterized by the fixed-point problem $(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})=f(\hat{\theta}_{C_{j}},\hat{\nu}_{C_{j}})$, where $f$ is as defined in (16). Iterative fixed-point methods (Picard, 1890; Lindelöf, 1894; Banach, 1922; Ryu & Boyd, 2016; Yang et al., 2021) solve such problems by considering an update rule of the form
@@
The form of (17) inspires us to approximate the true child solution $\hat{\theta}_{C_{j}}$ using a single fixed-point update taken from the parent solution $\hat{\theta}_{P}$:
@@
$=$ $\hat{\theta}_{P}-\frac{\eta}{n_{C_{j}}}\xi^{\top}\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}),$ (18)
where the product with $\xi^{\top}$ is interpreted similarly to its role in the gradient-based approximation (10) and to express the update (17) solely in terms of the target $\theta$-quantity. We interpret $\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}$ as an approximation of $\hat{\theta}_{C_{j}}$ obtained by taking a step from $\hat{\theta}_{P}$ in a direction that reduces the magnitude of the local estimating function $\Psi_{C_{j}}$. Notably, the approximation $\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ does not involve the $A_{P}$ matrix, relying only on the scores $\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$ evaluated at the parent solutions. In general, removing the inverse $A_{P}^{-1}$ provides computational cost savings of $\mathcal{O}(K^{3})$. The corresponding splitting criterion, which uses the fixed-point approximations $\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ as substitutes for $\hat{\theta}_{C_{j}}$ is given by
@@
Revisiting the varying coefficient example from Section 3.2, we see that splits based on fixed-point approximations $\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ are significantly more stable than those based on $\tilde{\theta}_{C_{j}}^{\texttt{grad}}$. Specifically, Figure 1 illustrates that splits that maximize $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ are more robust to ill-conditioning in the underlying local Jacobian $\nabla_{(\theta,\nu)}\mathbb{E}[\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})\mid X_{i}\in P]$, as is the case for highly correlated regressors in the varying coefficient model (14), and leading to highly stable splits.